# Differential Privacy of Cross-Attention with Provable Guarantee

### v2

@@ G.1 Privacy and Accuracy Analysis for High Dimensional Weighted Distance
3:     ${\mathcal{D}}_{1},\dots,{\mathcal{D}}_{d}:\textsc{DPTreeDistance}$triangleright Alg. 5
@@
7:procedure Init($X\subset[0,R]^{d}$, $n\in\mathbb{N}_{+}$, $w\in[-R_{w},R_{w}]^{n}$, $\epsilon\in(0,1)$, $\delta\in(0,1)$, $\delta^{\prime}\in(0,1)$, $c\in(0,0.1)$) $\triangleright$ Lemma E.2
@@
11:         $\mathcal{D}_{i}.$Init($X_{:,i}$, $n$, $w$, $c\epsilon/\sqrt{d\log(1/\delta^{\prime})}$, $\delta/d$) $\triangleright$ Alg. 5
@@
14:procedure DistanceQuery($y\in[0,R]^{d}$, $\alpha\in(0,1)$) $\triangleright$ Lemma G.1, Lemma G.2
@@
17:         Value $\leftarrow$ Value + $\mathcal{D}_{i}$.DistanceQuery$(y_{i},\alpha)$triangleright Alg. 6
@@
Lemma G.1 (Privacy of DistanceQuery, Algorithm 7).
@@
  • Let $A_{*}=\sum_{i\in[n]}w_{i}\|y-x_{i}\|_{1}$ represent the true distance query value.
@@
The $(\epsilon,\delta+\delta^{\prime})$-DP guarantee follows from the approximate DP advanced composition result Theorem B.8. Our algorithm instantiate each one-dimensional data structure with $(c\epsilon/\sqrt{d\log(1/\delta^{\prime})},\delta/d)$-DP total $d$ times.
From advanced composition in Theorem B.8, for a sufficient small parameter $\epsilon$ and constant $c$, we have the final privacy loss parameter be:
@@
Lemma G.2 (Accuracy of DistanceQuery, Algorithm 7).
@@
  • Let $A_{*}=\sum_{i\in[n]}w_{i}\|y-x_{i}\|_{1}$ represent the true distance query value.
@@
$|A-A_{*}|\leq\alpha A_{*}+O(\epsilon^{-1}\alpha^{-1/2}RR_{w}d\sqrt{\log(1/\delta^{\prime})}\cdot\log^{3/2}n).$
@@
We follow the similar idea in the proof of Lemma F.5. Let $z_{j,i}$ be the random variables that represent $z_{j}$ (used in the proof of Lemma F.5) for the $i$-th coordinate. We can observe that the overall error across $d$ coordinates can be upper bounded by
$|\sum_{i=1}^{d}\sum_{j=0}^{O(\log(n)/\alpha)}\frac{Rz_{j,i}}{(1+\alpha)^{j}}|$
where each $z_{j,i}$ is the sum of $O(\log n)$ truncated Laplace random variables independent to others. With $\epsilon$ scaled down by $c\epsilon/\sqrt{d\log(1/\delta^{\prime})}$ and $\delta$ scaled down by $\delta/d$, the variance of each individual dimension is given by (see Eq. (5))
$O(\alpha^{-1}\epsilon^{-2}dR^{2}R_{w}^{2}\log(1/\delta^{\prime})\log^{3}n).$
@@
$O(\alpha^{-1}\epsilon^{-2}d^{2}R^{2}R_{w}^{2}\log(1/\delta^{\prime})\log^{3}n).$
Finally, from Lemma B.3, we have the additive error given by
$O(\alpha^{-1/2}\epsilon^{-1}dRR_{w}\sqrt{\log(1/\delta^{\prime})}\cdot\log^{3/2}n).$
G.2 High Dimension Single Data Structure
We have the data structure that can solve weighted $\ell_{1}$-distance problem in $d$-dimensional data.
Theorem G.3 (DPTreeHighDim data structure).
There is a data structure DPTreeHighDim (Algorithm 7) that uses $O(nd)$ spaces to solve weighted $\ell_{1}$-distance query problem for dataset $X\subset[0,R]^{d}$ and support the following operations:
  • Init$(X\subset[0,R]^{d},n\in\mathbb{N}_{+},w\in[-R_{w},R_{w}]^{n},\epsilon\in(0,1),\delta\in(0,1),\delta^{\prime}\in(0,1),c\in(0,0.1))$. (Algorithm 7) It takes $O(nd)$ time to initialize the data structure.
  • DistanceQuery$(y\in[0,R]^{d},\alpha\in(0,1))$. (Algorithm 7) It takes $O(\alpha^{-1}d\log^{2}n)$ time to output a number $z$ such that
    • the process of output $z$ satisfies is $(\epsilon,\delta+\delta^{\prime})$-DP private, which computes $\sum_{i\in[n]}w_{i}\|y-x_{i}\|_{1}$,
    • $|z-\sum_{i\in[n]}w_{i}\|y-x_{i}\|_{1}|\leq\alpha\sum_{i\in[n]}w_{i}\|y-x_{i}\|_{1}+O(\epsilon^{-1}\alpha^{-1/2}RR_{w}d\sqrt{\log(1/\delta^{\prime})}\cdot\log^{3/2}n)$,
@@ G.2 High Dimension Single Data Structure
For the runtime analysis, since we loop data structure DPTreeDistance $d$ times, an additional $d$ factor will appear for both initialization and query time complexity. The DP is proved by Lemma G.1. The accuracy is proved by Lemma G.2. ∎
Appendix H Adaptive Query
In this section, we introduce how we can solve the adaptive query problem by our algorithm, using some tools from QRS+ [86]. Our idea is that, if we can prove that our algorithm can solve any query in the query space with certain error. Then, since adaptive query must lie in this space, we can handle adaptive query. In Section H.1, we show how we can boost the constant probability of our algorithm to high probability. In Section H.2, we show how we can apply the notion of $\epsilon_{0}$-net and bound all query points in net. In Section H.3, we show how we can bound all points in the query space by introducing an additive error. In Section H.4, we examine the effects of different norms on our adaptive query proof.
First, from Theorem G.3, given query $y\in[0,R]^{d},\alpha\in(0,1)$ we have DistanceQuery$(y,\alpha)$ that can solve $d$-dimension weighted $\ell_{1}$-distance problem with constant probability $0.99$. Now we show how to improve it to solve adaptive query problem.
H.1 Boost the Constant Probability to High Probability
We can repeat the data structure multiple times and take the median to boost the constant probability using Chernoff bound from Lemma B.2.
Lemma H.1 (Using Chernoff bound to boost the probability).
@@ H.1 Boost the Constant Probability to High Probability
  • Let relative error parameter $\alpha\in(0,1)$, the failure probability $p_{f}\in(0,0.01)$.
@@
  • Let $A_{*}=\sum_{i\in[n]}w_{i}\|y-x_{i}\|_{1}$ be the true answer.
  • Let $B=O(\epsilon^{-1}\alpha^{-1/2}lRR_{w}d\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$.
Then for each fixed query point $y$, we can have the process of outputting the median of $l$ responses is $(\epsilon,\delta+\delta^{\prime})$-DP and the error is upper bounded by $\alpha A_{*}+B$ with probability $1-p_{f}$.
@@
By basic composition Fact B.6, we prove the DP. Similar to the proof of Theorem D.2, we prove the error by Chernoff bound (Lemma B.2). ∎
H.2 From Each Fixed Query Point to All On-net Points
@@
Definition H.2 ($\ell_{p}$ $\epsilon_{0}$-net, see Definition 4.2.1 in Ver [96]).
We define $N$ be $\ell_{p}$ $\epsilon_{0}$-net of :=${\mathcal{B}}:=\{q\in[0,R]^{d}\}$ such that, for every point $q$ in ${\mathcal{B}}$, there exists $y\in N$ satisfying $\|y-q\|_{p}\leq\epsilon_{0}$.
Fact H.3 ($\ell_{\infty}$ $\epsilon_{0}$-net).
@@
Fact H.4 ($\ell_{2}$ $\epsilon_{0}$-net, see Lemma 5 in Woo [102]).
@@
Fact H.5 ($\ell_{1}$ $\epsilon_{0}$-net, see Theorem 2 in GS [41]).
@@
Lemma H.6 (From for each query point to for all points in net).
@@
  • Let relative error parameter $\alpha\in(0,1)$, the failure probability $p_{f}\in(0,0.01)$.
@@
  • Let $A_{*}=\sum_{i\in[n]}w_{i}\|y-x_{i}\|_{1}$ be the true answer.
  • Let $B=O(\epsilon^{-1}\alpha^{-1/2}lRR_{w}d\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$.
Then with probability $1-p_{f}$, for all query points $y\in N$, we can have the process of outputting the median of $l$ responses is $(\epsilon,\delta+\delta^{\prime})$-DP and the error is upper bounded by $\alpha A_{*}+B$.
@@
By basic composition Fact B.6, we prove the DP. From Lemma H.1, we know for each $y\in N$, the error is upper bounded by $\alpha A_{*}+B$ with probability $1-p_{f}/|N|$.
Then, by union bound, with probability $1-p_{f}$, the error of all $|N|$ query points in the net $y\in N$ is upper bounded by $\alpha A_{*}+B$. ∎
H.3 From Net Points to All Points
@@
Lemma H.7 (Lipschitz of query function).
@@
Lemma H.8 (From points in net to all points in query space).
@@
  • Let relative error parameter $\alpha\in(0,1)$, the failure probability $p_{f}\in(0,0.01)$.
@@
  • Let $f(y):=\mathrm{Median}(\{\textsc{DPTreeHighDim}_{j}.\textsc{DistanceQuery}(y,\alpha)\}_{j=1}^{l})$.
@@
  • Let $B=O(\epsilon^{-1}\alpha^{-1/2}lRR_{w}d\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$.
@@
By basic composition Fact B.6, we prove the DP.
@@
$~{}|f(y)-Z(y)|\leq\alpha Z(y)+B.$
From Lemma H.1, with $l=O(\log(|N|/p_{f}))$ we know
@@
where the first step follows from definition of $l$, the second step follows from Fact H.3.
We condition on event $E$ to be held. Then, by definition of $\ell_{\infty}$ $\epsilon_{0}$-net (see Definition H.2), for each $q\notin N$, there exists $y\in N$ such that
@@
where the first step follows from Lemma H.7, the second step follows from $\|x\|_{1}\leq d\|x\|_{\infty}$ for $x\in\mathbb{R}^{d}$, and the last step follows from Eq. (6).
@@
$\leq$ $~{}\alpha Z(y)+B+L\cdot d\cdot\epsilon_{0}$
$\leq$ $~{}\alpha Z(q)+B+2L\cdot d\cdot\epsilon_{0}$ (8)
where the first step follows from triangular inequality, the second step follows from Eq. (H.3), the third step follows from Lemma H.6, and the last step follows from Eq. (H.3).
@@
H.4 Effect of Different Norms on the Result
In the above proof, we have two different measure spaces, i.e. $\ell_{\infty}$ distance of $\epsilon_{0}$-net (Definition H.2) and $\ell_{1}$ Lipschitz (Lemma H.7).
One might ask, will the norm we choose in two spaces have an impact on the final result? We can show that the norm we choose currently is sufficient to use.
For different norms, the only differences in the proofs will be Lipschitz smoothness in Eq. (H.3) and the cardinality of $\epsilon_{0}$-net, i.e. $|N|$ in Fact H.3.
Lemma H.9.
If we use $\ell_{\infty}$ $\epsilon_{0}$-net and use $\ell_{1}$ Lipschitz in Lemma H.8, we have copies of data structure $l=O(d\log(nR/p_{f}))$.
Proof.
If we use $\ell_{\infty}$ to bound the distance to net, Eq. (H.3) is:
$|Z(y)-Z(q)|\leq$ $~{}nR_{w}\cdot\|y-q\|_{1}$
$\leq$ $~{}nR_{w}\cdot d\|y-q\|_{\infty}$
$\leq$ $~{}nR_{w}\cdot d\epsilon_{0}$
where the first step follows from Lemma H.7, the second step follows from $\|x\|_{1}\leq d\|x\|_{\infty}$ for $x\in\mathbb{R}^{d}$, and the last step follows from $\ell_{\infty}$ $\epsilon_{0}$-net.
Then, Eq. (H.3) is
$|f(y)-Z(q)|\leq$ $~{}\alpha Z(q)+B+2nR_{w}\cdot d\cdot\epsilon_{0}$
For $\ell_{\infty}$ distance, we have $|N|\leq(5R/\epsilon_{0})^{d}$ in Fact H.3.
We can choose $\epsilon_{0}=\Theta(1/n)$ to hide $nR_{w}\cdot d\cdot\epsilon_{0}$ term in $B$ in Lemma H.8. Thus,
$l=$ $~{}O(\log(|N|/p_{f})$
$=$ $~{}O(\log((R/\epsilon_{0})^{d}/p_{f})$
$=$ $~{}O(\log((nR)^{d}/p_{f}))$
$=$ $~{}O(d\log(nR/p_{f}))$
where the last step follows from $\log(a^{d}/b)=O(d\log(a/b))$ for any $a>1,0<b<1,d>1$. ∎
Lemma H.10.
If we use $\ell_{2}$ $\epsilon_{0}$-net and use $\ell_{1}$ Lipschitz in Lemma H.8, we have copies of data structure $l=O(d\log(nR/p_{f}))$.
Proof.
If we use $\ell_{2}$ to bound the distance to net, Eq. (H.3) changes to be:
$|Z(y)-Z(q)|\leq$ $~{}nR_{w}\cdot\|y-q\|_{1}$
$\leq$ $~{}nR_{w}\cdot\sqrt{d}\cdot\|y-q\|_{2}$
$\leq$ $~{}nR_{w}\cdot\epsilon_{0}\sqrt{d}$
where the first step follows from Lemma H.7, the second step follows from $\|x\|_{1}\leq\sqrt{d}\cdot\|x\|_{2}$ for $x\in\mathbb{R}^{d}$, and the last step follows from $\ell_{2}$ $\epsilon_{0}$-net.
Then, Eq. (H.3) changes to be
$|f(y)-Z(q)|\leq$ $~{}\alpha Z(q)+B+2nR_{w}\cdot\epsilon_{0}\sqrt{d}$
For $\ell_{2}$ distance, we also have $|N|\leq(5R/\epsilon_{0})^{d}$ in Fact H.4.
We can choose $\epsilon_{0}=\Theta(1/n)$ to hide $nR_{w}\cdot\sqrt{d}\cdot\epsilon_{0}$ term in $B$ in Lemma H.8. Thus,
$l=$ $~{}O(\log(|N|/p_{f})$
$=$ $~{}O(\log((R/\epsilon_{0})^{d}/p_{f})$
$=$ $~{}O(\log((nR)^{d}/p_{f}))$
$=$ $~{}O(d\log(nR/p_{f}))$
where the last step follows from $\log(a^{d}/b)=O(d\log(a/b))$ for any $a>1,0<b<1,d>1$. ∎
Lemma H.11.
If we use $\ell_{1}$ $\epsilon_{0}$-net and use $\ell_{1}$ Lipschitz in Lemma H.8, we have copies of data structure $l=O(d\log(ndR/p_{f}))$.
Proof.
If we use $\ell_{1}$ to bound the distance to net, Eq. (H.3) changes to be:
$|Z(y)-Z(q)|\leq$ $~{}nR_{w}\cdot\|y-q\|_{1}$
$\leq$ $~{}nR_{w}\cdot\epsilon_{0}$
where the first step follows from Lemma H.7, and the last step follows from $\ell_{1}$ $\epsilon_{0}$-net.
Then, Eq. (H.3) changes to be
$|f(y)-Z(q)|\leq$ $~{}\alpha Z(q)+B+2nR_{w}\cdot\epsilon_{0}$
For $\ell_{1}$ distance, we have $|N|\leq(5R\sqrt{d}/\epsilon_{0})^{d}$.
We can choose $\epsilon_{0}=\Theta(1/n)$ to hide $nR_{w}\cdot\epsilon_{0}$ term in $B$ in Lemma H.8. Thus,
$l=$ $~{}O(\log(|N|/p_{f})$
$=$ $~{}O(\log((R\sqrt{d}/\epsilon_{0})^{d}/p_{f})$
$=$ $~{}O(\log((nR\sqrt{d})^{d}/p_{f}))$
$=$ $~{}O(d\log(nRd/p_{f}))$
where the last step follows from $\log(a^{d}/b)=O(d\log(a/b))$ for any $a>1,0<b<1,d>1$. ∎
From the above analysis, we can show that $\ell_{\infty}$ or $\ell_{2}$ $\epsilon_{0}$-net is slightly better than $\ell_{1}$ $\epsilon_{0}$-net.
  • $\ell_{\infty}$ $\epsilon_{0}$-net, Lemma H.9: we have $l=O(d\log(nR/p_{f}))$.
  • $\ell_{2}$ $\epsilon_{0}$-net, Lemma H.10: we have $l=O(d\log(nR/p_{f}))$.
  • $\ell_{1}$ $\epsilon_{0}$-net, Lemma H.11: we have $l=O(d\log(nRd/p_{f}))$.
Thus, the norm we choose for $\epsilon_{0}$-net is sufficient good.
Appendix I Softmax Activation
In this section, we introduce how we extend previous $\ell_{1}$ distance results to the Softmax activation function, which is the most widely used distance measure in attention mechanism based models.
In Section I.1, we show how to extend to the Softmax distance function in Lemma I.6. In Section I.2, we show how to adjust our algorithms. In Section I.3, we extend our algorithm to be robust to adaptive query. In Section I.4, we give the proof of our main result Theorem 3.1.
I.1 Exponential Inner Product
In this section, we show how we obtain the Softmax distance using $\ell_{2}^{2}$ distance query. First, we provide some helpful results from AS [6].
Definition I.1 (Definition 3.1 in AS [6]).

### v3

@@ E.1 Privacy and Accuracy Analysis for High Dimensional Weighted Distance
3:   ${\mathcal{D}}_{1},\dots,{\mathcal{D}}_{d}:\textsc{DPTreeDistance}$ $\triangleright$ Alg. 4
@@
7:procedure Init($X\subset[0,R]^{d}$, $n\in\mathbb{N}_{+}$, $w\in[-R_{w},R_{w}]^{n}$, $\epsilon\in(0,1)$, $\delta\in(0,1)$, $\delta^{\prime}\in(0,1)$, $c\in(0,0.1)$)
@@
11:   $\mathcal{D}_{i}.$Init($X_{:,i}$, $n$, $w$, $c\epsilon/\sqrt{d\log(1/\delta^{\prime})}$, $\delta/d$) $\triangleright$ Alg. 4
@@
14:procedure DistanceQuery($y\in[0,R]^{d}$) $\triangleright$ Lemma E.1, Lemma E.2
@@
[…]
@@
Lemma E.1 (Privacy of DistanceQuery, Algorithm 6).
@@
  • Let $A_{*}=\sum_{i\in[n]}w_{i}\|y-x_{i}\|_{p}^{p}$ represent the true distance query value.
@@
The $(\epsilon,\delta+\delta^{\prime})$-DP guarantee follows from the approximate DP advanced composition result Theorem A.10. Our algorithm instantiate each one-dimensional data structure with $(c\epsilon/\sqrt{d\log(1/\delta^{\prime})},\delta/d)$-DP total $d$ times.
From advanced composition in Theorem A.10, for a sufficient small parameter $\epsilon$ and constant $c$, we have the final privacy loss parameter be:
@@
Lemma E.2 (Accuracy of DistanceQuery, Algorithm 6).
@@
  • Let $A_{*}=\sum_{i\in[n]}w_{i}\|y-x_{i}\|_{p}^{p}$ represent the true distance query value.
@@
$|A-A_{*}|\leq O(\epsilon^{-1}dp(2R)^{p}R_{w}\sqrt{\log(1/\delta^{\prime})}\cdot\log^{3/2}n).$
@@
We follow the similar idea in the proof of Theorem D.1. With $\epsilon$ scaled down by $c\epsilon/\sqrt{d\log(1/\delta^{\prime})}$ and $\delta$ scaled down by $\delta/d$, the variance of each individual dimension is given by (see proof of Theorem D.1)
$O(\epsilon^{-2}dp^{2}(2R)^{2p}R_{w}^{2}\log(1/\delta^{\prime})\log^{3}n).$
@@
$O(\epsilon^{-2}d^{2}p^{2}(2R)^{2p}R_{w}^{2}\log(1/\delta^{\prime})\log^{3}n).$
Finally, from Lemma A.3, we have the additive error given by
$O(\epsilon^{-1}dp(2R)^{p}R_{w}\sqrt{\log(1/\delta^{\prime})}\cdot\log^{3/2}n).$
E.2 High Dimension Single Data Structure
We have the data structure that can solve weighted $\ell_{p}^{p}$ distance problem in $d$-dimensional data.
Theorem E.3 (DPTreeHighDim data structure).
There is a data structure DPTreeHighDim (Algorithm 6) that uses $O(npd)$ spaces to solve weighted $\ell_{p}^{p}$ distance query problem for dataset $X\subset[0,R]^{d}$ and support the following operations:
[…]
    • the process of output $z$ satisfies is $(\epsilon,\delta+\delta^{\prime})$-DP private, which computes $\sum_{i\in[n]}w_{i}\|y-x_{i}\|_{p}^{p}$,
    • $|z-\sum_{i\in[n]}w_{i}\|y-x_{i}\|_{1}|\leq O(\epsilon^{-1}dp(2R)^{p}R_{w}\sqrt{\log(1/\delta^{\prime})}\cdot\log^{3/2}n)$,
@@ E.2 High Dimension Single Data Structure
For the runtime analysis, since we loop data structure DPTreeDistance $d$ times, an additional $d$ factor will appear for both initialization and query time complexity. The DP is proved by Lemma E.1. The accuracy is proved by Lemma E.2. ∎
Appendix F Adaptive Query
In this section, we introduce how we can solve the adaptive query problem by our algorithm, using some tools from [70]. Our idea is that, if we can prove that our algorithm can solve any query in the query space with certain error. Then, since adaptive query must lie in this space, we can handle adaptive query. In Section F.1, we show how we can boost the constant probability of our algorithm to high probability. In Section F.2, we show how we can apply the notion of $\epsilon_{0}$-net and bound all query points in net. In Section F.3, we show how we can bound all points in the query space by introducing an additive error.
First, from Theorem E.3, given query $y\in[0,R]^{d}$ we have DistanceQuery(y)(y) that can solve $d$-dimension weighted $\ell_{p}^{p}$ distance problem with constant probability $0.99$. Now we show how to improve it to solve adaptive query problem. Here, we focus on the case when $p=1$.
F.1 Boost the Constant Probability to High Probability
We can repeat the data structure multiple times and take the median to boost the constant probability using Chernoff bound from Lemma A.2.
Lemma F.1 (Using Chernoff bound to boost the probability).
@@ F.1 Boost the Constant Probability to High Probability
  • Let the failure probability $p_{f}\in(0,0.01)$.
@@
  • Let $B=O(\epsilon^{-1}lRR_{w}d\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$.
Then for each fixed query point $y$, we can have the process of outputting the median of $l$ responses is $(\epsilon,\delta+\delta^{\prime})$-DP and the error is upper bounded by $B$ with probability $1-p_{f}$.
@@
By basic composition Fact A.8, we prove the DP. Similar to the proof of Theorem B.2, we prove the error by Chernoff bound (Lemma A.2). ∎
F.2 From Each Fixed Query Point to All On-net Points
@@
Definition F.2 ($\ell_{p}$ $\epsilon_{0}$-net, see Definition 4.2.1 in [77]).
We define $N$ be $\ell_{p}$ $\epsilon_{0}$-net of ${\mathcal{B}}:=\{q\in[0,R]^{d}\}$ such that, for every point $q$ in $\mathcal{B}$}, there exists $y\in N$ satisfying $\|y-q\|_{p}\leq\epsilon_{0}$.
Fact F.3 ($\ell_{\infty}$ $\epsilon_{0}$-net).
@@
Fact F.4 ($\ell_{2}$ $\epsilon_{0}$-net, see Lemma 5 in [83]).
@@
Fact F.5 ($\ell_{1}$ $\epsilon_{0}$-net, see Theorem 2 in [28]).
@@
Lemma F.6 (From for each query point to for all points in net).
@@
  • Let the failure probability $p_{f}\in(0,0.01)$.
@@
  • Let $B=O(\epsilon^{-1}lRR_{w}d\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$.
Then with probability $1-p_{f}$, for all query points $y\in N$, we can have the process of outputting the median of $l$ responses is $(\epsilon,\delta+\delta^{\prime})$-DP and the error is upper bounded by $B$.
@@
By basic composition Fact A.8, we prove the DP. From Lemma F.1, we know for each $y\in N$, the error is upper bounded by $B$ with probability $1-p_{f}/|N|$.
Then, by union bound, with probability $1-p_{f}$, the error of all $|N|$ query points in the net $y\in N$ is upper bounded by $B$. ∎
F.3 From Net Points to All Points
@@
Lemma F.7 (Lipschitz of query function).
@@
Lemma F.8 (From points in net to all points in query space).
@@
  • Let the failure probability $p_{f}\in(0,0.01)$.
@@
  • Let $f(y):=\mathrm{Median}(\{\textsc{DPTreeHighDim}_{j}.\textsc{DistanceQuery}(y)\}_{j=1}^{l})$.
@@
  • Let $B=O(\epsilon^{-1}lRR_{w}d\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$.
@@
By basic composition Fact A.8, we prove the DP.
@@
$~|f(y)-Z(y)|\leq B.$
From Lemma F.1, with $l=O(\log(|N|/p_{f}))$ we know
@@
where the first step follows from definition of $l$, the second step follows from Fact F.3.
We condition on event $E$ to be held. Then, by definition of $\ell_{\infty}$ $\epsilon_{0}$-net (see Definition F.2), for each $q\notin N$, there exists $y\in N$ such that
@@
where the first step follows from Lemma F.7, the second step follows from $\|x\|_{1}\leq d\|x\|_{\infty}$ for $x\in\mathbb{R}^{d}$, and the last step follows from Eq. (6).
@@
$\leq$ $~B+L\cdot d\cdot\epsilon_{0}$ (8)
where the first step follows from triangular inequality, the second step follows from Eq. (F.3), the third step follows from Lemma F.6.
@@
Appendix G Softmax Activation
In this section, we introduce how we extend previous $\ell_{p}^{p}$ distance results to the Softmax activation function, which is the most widely used distance measure in attention mechanism based models.
In Section G.1, we show how to extend to the Softmax distance function in Lemma G.6. In Section G.2, we show how to adjust our algorithms. In Section G.3, we extend our algorithm to be robust to adaptive query. In Section G.4, we give the proof of our main result Theorem 4.1.
G.1 Exponential Inner Product
In this section, we show how we obtain the Softmax distance using $\ell_{2}^{2}$ distance query. First, we provide some helpful results from [4].
Definition G.1 (Definition 3.1 in [4]).